OpenStudy (anonymous):

find 3Sin^2x-2Cos^2x-Sinx=0

OpenStudy (dumbcow):

sin^2 + cos^2 = 1 ----> cos^2 = 1 - sin^2 \[3 \sin^2 x -2(1-\sin^2 x) -\sin x = 0\] \[5 \sin^2 x - \sin x - 2 = 0\] solve it like a quadratic u = sin x \[5u^2 - u -2 = 0\]

OpenStudy (anonymous):

Tyvm